OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. 3 – 7x + 2x when x = 3 A. –30 B. –12 C. –18 D. –24

OpenStudy (anonymous):

it is going to look like this right: 3 - 7 * 3 + 2 * 3

OpenStudy (anonymous):

i think so... idk my math grades went down hill after 5th grade so i have no idea

OpenStudy (anonymous):

what grade r u in?

OpenStudy (anonymous):

7th...

OpenStudy (anonymous):

i am in 7th grade and math is my best subject...

OpenStudy (anonymous):

oh COOL!!

OpenStudy (anonymous):

Im a nerd in ALL other classes tho its just when the alphabet became a grade i paniced

OpenStudy (anonymous):

try evaluating what i told u just TRY

OpenStudy (anonymous):

I just need the answer tho

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

ohh dont worry u will get better trust me.. just like me in 4th grade i was always getting f and d's i promise..

OpenStudy (anonymous):

just try

OpenStudy (anonymous):

umm 15

OpenStudy (anonymous):

I HAVE NO CLEAR IDEA MY MIND IS FOGGED

OpenStudy (anonymous):

let me try it..

OpenStudy (anonymous):

remember PEMDAS

OpenStudy (anonymous):

sorry caps

OpenStudy (anonymous):

follow that order

OpenStudy (anonymous):

what on this great planet is that?

OpenStudy (anonymous):

P Parentheses first E Exponents (ie Powers and Square Roots, etc.) MD Multiplication and Division (left-to-right) AS Addition and Subtraction (left-to-right) WOAH U NEVER LEARNED IT WOW!!

OpenStudy (anonymous):

i WAS smart in gt to its math kicks me un the hind end...

OpenStudy (anonymous):

3-7(3)+2(3) PEMDAS -7 x 3 = -21 3-21+2(3) 2 x 3 = 6 3-21+6 = -12 Your answer is B

OpenStudy (anonymous):

yup u r correct bc a negative * a Positive is A Negative

OpenStudy (anonymous):

thank you soo much

OpenStudy (anonymous):

sorry got mad...

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

pig presedent i have no idea what u said about please excuse my dear aunt sally

OpenStudy (anonymous):

Evaluate the expression, using the given value of the variable. z ÷ 2 – 4z + 3(z – 3) when z = 8 A. –17 B. –13 C. –7 D. –1 this one :D

OpenStudy (anonymous):

just look at this site http://www.mathsisfun.com/operation-order-pemdas.html

OpenStudy (anonymous):

i cant other sights are blocked.

OpenStudy (anonymous):

oh like me too...my parents r like that

OpenStudy (anonymous):

but this is school related

OpenStudy (anonymous):

cant mom doesnt want me 2

OpenStudy (anonymous):

ok